. Let B be the number of boys in a family of 4 children. Construct a probability distribution of the random variable x. Construct a probability histogram 


1
Expert's answer
2022-03-24T05:44:22-0400

Let probability of boy or girl is equal 0.5. 1 boy from 4 or 1 girl from 4 we can choose for 4 ways ("C_4^1=\frac{4!}{1!3!}=4)" . And 2 boys from 4 we can choose for 6 ways ("C_4^2=\frac{4!}{2!2!}=6" .

p- probability of boy ,equal 0.5

q - probability of girl, equal 0.5

"P(0)=q^4=0.5^4=0.0625"

"P(1)=4q^3p=4(0.5)^4=0.25"

"P(2)=6p^2q^2=6(0.5)^4=0.375"

"P(3)=4pq^3=4(0.5)^4=0.25"

"P(4)=p^4=0.0625"